Keep it​ Minimal, Inc. is a manufacturer of furniture. At the beginning of​ February, it was estimated that each unit of Let the Light In would require 4 hours of direct​ labor, and the related direct labor cost was expected to equal​ $36 per unit of product. During​ February, the company logged​ 3,750 hours of direct labor hours to produce​ 1,000 units of Let the Light In. Direct laborers were paid at a rate of​ $9.15 per hour. Which of the following statements is correct with regard to Let the Light​ In’s February​ production?

A. The total actual direct labor cost was higher than the total expected direct labor cost.

B. It is likely that the direct labor variances are not related since both the rate variance and the efficiency variance are unfavorable.

C. If the company used higher paid and more efficient workers than expected for production during the​ period, it was not​ “worth ​it.”

D. The direct labor hours expected for the actual output during the month were more than the actual direct labor hours logged during the month.

E. The direct labor standard rate was greater than the direct labor actual rate during the month.
Answer #1
Standard direct labor rate per hour = 36/4 = $9.00
Standard direct labor hours for actual production = 1000*4 = 4000
Total expected direct labor cost = 4000*9 = $36000
Total actual direct labor cost = 3750*9.15= $34312
The direct labor hours expected for the actual output during the month were more than the actual direct labor hours logged during the month.
Option D is correct